The half-value of radium is 1590 years. in how many years will one grams of pure element lose one centigram
Answer by Boreal(15235)

A centigram is (1/100) of a gram
(1/2)^n=0.99; n log(1/2)=log (0.99)
-n log (2)=log (0.99)
-0.30103n=-0.00436
n=0.00436/0.30103
n=0.1448 half lives
0.0145*1590=23.03 years
-------------------------